OpenStudy (anonymous):

PLEASE HELP ME! (Logarithm Question) Prove that:

OpenStudy (anonymous):

\[\frac{ 1 }{ \log _{x}10 } + \frac{ 1 }{ \log _{y}10} = \frac{ 1 }{ \log _{x _{y}}10}\]

OpenStudy (anonymous):

add on the left, then use the change of base formula, if i remember correctly seems like this question comes up every month or so

OpenStudy (anonymous):

hi @satellite73 !! since you've helped me so much before, when you are done here could you possibly help me with my recent question?

OpenStudy (anonymous):

I`m sorry but I am totally clueless on this question :S Can you show me how to do it plz

OpenStudy (anonymous):

this is going to be a bear to write add on the left you get \[\frac{\log_x(10)+\log_y(10)}{\log_x(10)\times \log_y(10)}\]now change of base. doesn't matter what base you use, so you might as well use base ten \[\frac{\frac{\log(10)}{\log_x(10)}+\frac{\log(10)}{\log_y(10)}}{\frac{\log(10)}{\log_x(10)}\times \frac{\log(10)}{\log_y(10)}}\]

OpenStudy (anonymous):

now since \(\log(10)=1\) you can rewrite all this mess as \[\frac{\frac{1}{\log_x(10)}+\frac{1}{\log_y(10)}}{\frac{1}{\log_x(10)}\times \frac{1}{\log_y(10)}}\]

OpenStudy (anonymous):

now add in the numerator, get \[\frac{\frac{\log_y(10)+\log_x(10)}{\log_x(10)\log_y(10)}}{\frac{1}{\log_x(10)\log_y(10)}}\]

OpenStudy (anonymous):

denominators cancel gives \[\log_x(10)+\log_y(10)\]

OpenStudy (anonymous):

would that be the final answerÉ

OpenStudy (anonymous):

no, we are not quite done we have to show the right hand side is the same thing

OpenStudy (anonymous):

all that was on the left. the right hand side is \[\frac{1}{\log_{xy}(10)}\] and again by change of base this is \[\frac{1}{\frac{\log(10)}{\log(xy)}}=\log(xy)=\log(x)+\log(y)\] oh damn i messed up somewhere

OpenStudy (anonymous):

ok it is much easier that i made it out to be the right hand side is \[\log(x)+\log(y)\] by what i wrote directly above, the change of base now lets do the same on the left

OpenStudy (anonymous):

\[\frac{1}{\log_x(10)}=\frac{1}{\frac{\log(10)}{\log(x)}}=\frac{1}{\frac{1}{\log(x)}}=\log(x)\] similarly \[\frac{1}{\log_y(10)}=\log(y)\]

OpenStudy (anonymous):

so both the left hand side and the right hand side turn in to \[\log(x)+\log(y)\]and i hope it is clear that when i write \(\log(x)\) i mean \(\log_{10}(x)\)

OpenStudy (anonymous):

yea its making sense

OpenStudy (anonymous):

ok it wasn't as hard as i made it out to be just the change of base formula on both sides to rewrite them both in base 10

OpenStudy (anonymous):

Thank you!

OpenStudy (dinnertable):

Thanks a lot for the help!